OpenStudy (anonymous):

what is the square root of -3 and 216?

OpenStudy (anonymous):

what does "and" mean in this context?

OpenStudy (anonymous):

sqrt ( -2+216) ?

OpenStudy (anonymous):

-3***

OpenStudy (anonymous):

\[\sqrt{216}=\sqrt{36\times 6}=\sqrt{36}\sqrt 6=6\sqrt6\]

OpenStudy (anonymous):

@satellite73 how did you enlargen your equation's fonts?

OpenStudy (anonymous):

none and 6 radical 6